Simplify the expressions or solve the equations (If you're solving the equation...STATE THE PROPERTIES)
a) 7x-4y-2y-9-4x+6
So far I have 7x-4x+4y-2y+(-9)+6
=3x+2y-3
=-6y
-combine like terms
-subtract
-solution
Is this correct or too far off, can you please give me the corrections if I am wrong, thank you.

Found 2 solutions by College Student, olya:
Answer by College Student(505)

You went one step too far.
You need to stop at =3x+2y-3 (...not sure where you got =-6y)
Everything else looks great. Good job!

Answer by olya(2)

hmm almost everything correct :
7x-4y-2y-9-4x+6=
(7x-4x)-(4y+2y)-(9-6)=
3x-6y-3